OpenStudy (anonymous):

Which property is demonstrated? 5 + 8 + 15 = 5 + 15 + 8 A. Associative Property of Addition B. Commutative Property of Multiplication C. Associative Property of Multiplication D. Commutative Property of Addition

OpenStudy (alexandervonhumboldt2):

Commutative Property of Addition means a + b = b + a.

OpenStudy (alexandervonhumboldt2):

associative property of addition, states that (a + b) + c = a + (b + c)